OpenStudy (kingj):

What is the sum of the arithmetic sequence 3, 9, 15…, if there are 34 terms? 2,467 2,684 3,072 3,468

OpenStudy (anonymous):

The sum of an arithmetic sequence is given by; S = n/2 [2(first) + (n-1)d], where d is the difference, 'first' is the first term, and n is the number of terms you are summing. Hence; S = 34/2 [2(3) + (33)6] = 3468